OpenStudy (anonymous):

A triangle with the sides 6', 8', and 10' is an example of the Pythagorean Theorem. True or False?

OpenStudy (anonymous):

A^2+B^2=C^2 a and b are the long and short legs of the triangle c is hypotenuse see if a+b=c